OpenStudy (anonymous):

Let T:R2 R2 be a linear transformation that sends the vector u=(5,2) into (2,1) and maps v=(1,3) into (−1,3). Use properties of a linear transformation to calculate 3U,2V and (3U+2V)

OpenStudy (anonymous):

Using property T(aU) = aT(U) of linear transformation: 3U = 3*(2,1) = (6,3) 2V = 2*(-1,3) = (-2,6) Using T(au+av) = aT(u) + bT(v) 3U+2V = (6,3) + (-2,6) = (4,9) Does this help ?

OpenStudy (anonymous):

thank you so much